# The ALMA Survey of Gas Evolution of PROtoplanetary Disks (AGE-PRO): I. Program Overview and Summary of First Results

### v1

@@ Abstract
We present the ALMA Survey of Gas Evolution of PROtoplanetary Disks (AGE-PRO), a Large Program of the Atacama Large Millimeter/submillimeter Array (ALMA). AGE-PRO aims to systematically trace the evolution of gas disk mass and size throughout the lifetime of protoplanetary disks. It uses a carefully selected sample of 30 disks around M3-K6 stars in three nearby star-forming regions: Ophiuchus (0.5-1 Myr), Lupus (1-3 Myr), and Upper Sco (2-6 Myr). Assuming the three regions had similar initial conditions and evolutionary paths, we find the median gas disk mass appears to decrease with age. Ophiuchus disks have the highest median gas mass (6 MJup), while the Lupus and Upper Sco disks have significantly lower median masses (0.68 and 0.44 MJup, respectively). Notably, the gas and dust disk masses appear to evolve on different timescales. This is evidenced by the median gas-to-dust mass ratio, which decreases from 122 in the youngest disks ($<$1 Myr) to 46 in Lupus disks, and then increases to 120 in the Upper Sco disks. The median gas disk sizes range between 74-110 au, suggesting that typical gas disks are much smaller than those of well-studied, massive disks. Population synthesis models suggest that magneto-hydrodynamic wind-driven accretion can reproduce median disk properties across all three regions, when assuming compact disks with a declining magnetic field over time. In contrast, turbulent-driven models overestimate gas masses of $>$1 Myr disks by an order of magnitude. Here we discuss the program’s motivation, survey design, sample selection, observation and data calibration processes, and highlight the initial results.
@@ 1 Introduction
Planets form in gas- and dust-rich circumstellar disks around young stars (e.g., see reviews by Williams & Cieza, 2011; Andrews, 2020; Öberg et al., 2023). Most of these disks last for up to 3-10 Myr before the gas and dust dissipate, which sets the timescale of giant planet formation (e.g., Fedele et al., 2010; Ribas et al., 2014). The structure and evolution of these disks profoundly affect every step of planet formation: from the initial grain growth, the formation of planetesimals, the accretion of planetary atmospheres, to the migration of planets (e.g., Benz et al., 2014; Drazkowska et al., 2023; Krijt et al., 2022).
@@ 2.1 Sample Selection
Figure 1: Top: Hertzsprung-Russell diagram of the AGE-PRO sample. The pre-main-sequence evolutionary tracks and isochrones are from Feiden (2016); Baraffe et al. (2015). Ophiuchus sources do not have well-constrained stellar luminosity due to high extinction. We instead plot their luminosity ranges based on statistical ages of Class I and Flat spectrum sources (0.5-1 Myr) (Evans et al., 2009). Bottom: 1.3mm-continuum luminosity ($L_{\rm mm}$) distributions (using AGE-PRO continuum fluxes and normalized to 140 pc) of the selected sample in each region. The sample covers the entire $L_{\rm mm}$ range of M3-K6 sources in each star-forming region (see detailed comparisons in Ruiz-Rodriguez et al., 2025, in press; Deng et al., 2025, in press; Agurto-Gangas et al., 2025, in press).
@@
References. — All distances are adopted from Bailer-Jones et al. (2021). In Col. (13), the references for Class, Spectral type, {$T_{\rm eff}$, $L_{\ast}$}, Av, and accretion rates: 1 = Evans et al. (2009), 2 = Cieza et al. (2019), 3 = Ruiz-Rodriguez et al. (2025, in press), 4 = Alcalá et al. (2014), 5 = Alcalá et al. (2017), 6 = Deng et al. (2025, in press), 7 = Luhman (2022), 8 = Manara et al. (2020), 9 = Carpenter et al. 2024, submitted, 10 = Manara et al. (2023), 11 = Fang et al. (2023), 12 = Agurto-Gangas et al. (2025, in press).
The AGE-PRO sample was selected with four main criteria: (1) sources with known stellar spectral type between M3-K6  [^1] Upper Sco 2 was classified as M2 in Preibisch et al. (2001) which we adopted in the AGE-PRO sample selection stage. But more recently Manara et al. (2020) reclassified the source as M4.5 with X-shooter spectra. , roughly corresponding to a stellar mass (M) of 0.3-0.8 M. The majority of stellar spectral types were adopted from literature and the rest from Ruiz-Rodriguez in prep. The literature stellar luminosities were updated with Gaia DR3 distances (see the list of references in Table 2.1). The focus on a narrow space of stellar mass is because disk properties are expected to depend on stellar mass (e.g., dust disk masses are known to be correlated with M; Pascucci et al. 2016; Andrews et al. 2018a). The stellar properties of the AGE-PRO sample are listed in Table 2.1. (2) sources without known companions or that are in wide-separation binaries ($>$600 AU), as close binaries may evolve differently due to tidal interactions (e.g., Cuello et al., 2023). (3) For the Ophiuchus region, Class I/Flat spectrum sources were chosen to ensure the sub-sample is $<$1 Myr old (Evans et al., 2009); for Lupus and Upper Sco regions, Class II sources were selected to exclude debris disks. (4) sources were selected from previous detections of mm continuum and CO line emission. These four criteria led to 10-34 viable candidates for each region. To ensure that our disks are representative of the selected M range, ten disks were selected from each region that covers the spread of continuum luminosities in the region (Ansdell et al. 2016; Barenfeld et al. 2016; Cieza et al. 2019; Carpenter et al. 2025). We excluded known edge-on disks ($>$70 deg) as they have smaller emitting areas and thus have fainter line emission. Figure 1 shows the stellar properties of the AGE-PRO sample in the Hertzsprung-Russell diagram and their continuum luminosity distributions.
@@ 2.2 Ages of the AGE-PRO sample
Since the primary goal of the AGE-PRO program is to study the evolution of gas disks, constraining the ages of the sample is crucial. However, estimating the absolute ages of very young stars ($<$10 Myr, all pre-main sequence PMS) is challenging due to the scarcity of direct age indicators and limited knowledge of star formation processes (see review by Soderblom et al. 2014). Typically, the ages of young PMSs are estimated by comparing observations with theoretical evolutionary tracks in either Hertzsprung-Russell Diagrams (HRDs) or color-magnitude diagrams (CMDs). If the stellar parameters and extinctions of individual stars are sufficiently constrained, ages can be determined fitting evolutionary tracks to each star, though these individual ages are often considered less reliable. Median ages are then adopted for the ages of clusters or star-forming regions. Another technique, often used when the extinctions and spectral types of individual stars are uncertain, is to derive a single age for an entire cluster by fitting all stars with one theoretical isochrone, assuming all clustered stars to be coeval.
@@ 2.3 Stellar masses of the AGE-PRO sample
In Figure 2 we compare the distribution of stellar masses across the three star-forming regions. The stellar masses are derived from the comparison with evolutionary tracks as described above. The bolometric temperature of our Ophiuchus sample indicates these are more evolved embedded disks, and therefore ,their current stellar masses should be close to their final masses. The median stellar masses of the three regions 0.45, 0.40, and 0.43 M for the Ophiuchus, Lupus, and Upper Sco region, respectively. The uncertainties of stellar masses from our estimations are typically 0.1-0.2 M, based on the uncertainties of the stellar effective temperature and luminosity (see details in Deng et al., 2025, in press; Agurto-Gangas et al., 2025, in press). However, these uncertainties do not include uncertainties of evolutionary tracks which are known to be model-dependent (Soderblom et al., 2014). Comparisons between stellar masses derived by evolutionary tracks with dynamical masses measured in binary systems suggest typical 30-80% uncertainties (Simon et al., 2019). Given these uncertainties on stellar masses, we consider the stellar masses in our sample are similar across the three regions and therefore suitable for comparison.
@@ 2.4.1 Lines to measure gas disk masses
But the conversion from CO gas mass to the total gas disk mass suffers from the high uncertainty of CO-to-H2 ratio ($x_{\rm CO}$) (Favre et al. 2013; Schwarz et al. 2016; Bergin & Williams 2017; Miotello et al. 2023). The uncertainty of $x_{\rm CO}$ likely increases over time due to chemical processes converting CO into more complex carbon carriers and dust growth carries CO ice from the disk atmosphere to the mid-plane (e.g, Schwarz et al., 2016; Bosman et al., 2018; Krijt et al., 2018, 2020). This expectation has been seen in some recent observations. Zhang et al. (2020); Bergner et al. (2020) compared CO isotopologue line fluxes with those predicted by thermo-chemical models and found that C/H and O/H elemental ratios of Class I/FS disks is consistent with the Interstellar Medium (ISM) ratio ($\sim$10-4). In contrast, they found that the C/H and O/H abundance of Class II disks needs to be reduced to $01\times$ 0.01 × ISM level. This is consistent with the timescales of the chemical processing of CO and the growth from ISM-sized dust particles to pebbles being $\sim$1 Myr (e.g., Schwarz et al., 2018; Bosman et al., 2018; Krijt et al., 2020). However, other works that include conversion of CO into CO2 ice do not require significant reductions in the C/H and O/H elemental abundances with respect to ISM values (Ruaud et al., 2022; Deng et al., 2023; Pascucci et al., 2023b; Ruaud & Gorti, 2024).
Fortunately, recent studies have provided a pathway forward to constrain $x_{\rm CO}$ and subsequently the gas disk mass. N2H+ line flux is sensitive to changes of $x_{\rm CO}$, because when gas-phase CO is around, N2H+ is hard to form, as CO gas competes with N2 for available H+${}_{3}^{+}$ that is needed for N2H+ formation. Furthermore, even if some N2H+ forms, it is rapidly destroyed by proton transfer to gas-phase CO. Therefore, the line flux ratios of CO isotopologue and N2H+ provide constraints of $x_{\rm CO}$ and thus largely improve the precision of gas mass estimates (Anderson et al. 2019, 2022; Trapman et al. 2022b; Sturm et al. 2023).
@@
Given these observations and theoretical expectations, the AGE-PRO program adopted different approaches to measure gas masses for sources younger and older than 1 Myr. For Ophiuchus sources ($<$1 Myr), we used C17O (2-1) and assume ISM level $x_{\rm CO}$ of 10-4 to measure gas disk masses. Based on the dust masses and an ISM level of CO abundance, radiative transfer models predict that C18O (2-1) is optically thick and therefore the more rare C17O line (typical C18O/C17O abundance ratio of 3.6, Wilson 1999) is used for mass measurements (Miotello et al., 2016). For Class II disks, $x_{\rm CO}$ needs to be constrained for accurate gas measurements, and therefore 13CO, C18O (2-1), and N2H+ (3-2) lines were used to constrain $x_{\rm CO}$ and the gas disk masses.
@@ 3 Observations
The ALMA observations of the AGE-PRO program were conducted with the 12-meter array from 2021 December to 2022 August, with baseline lengths between 15 to 2917 m. Observational logs of each execution, including observational dates, array configurations, baseline length range, number of antennas, phase and flux calibrators, are shown in Ruiz-Rodriguez et al. (2025, in press); Deng et al. (2025, in press); Agurto-Gangas et al. (2025, in press), for sources in Ophiuchus, Lupus, and Upper Sco, respectively.
[…]
@@ 3.1 Pipeline calibration and self-calibration
After the standard pipeline calibration, we conducted self-calibration using CASA version 6.4 to improve the signal-to-noise (SNR) of the data. Self-calibration of the continuum visibilities was performed for sources with continuum SNR larger than 20. Our methodology was based on the self-cal procedures of the DSHARP and MAPS large programs (Andrews et al., 2018b; Czekala et al., 2021), with modifications tailored to the AGE-PRO program. Unlike the targets of the DSHARP and MAPS large programs which were bright in the millimeter continuum, many of our sources only have intermediate and faint continuum emission ($<$10 mJy at 1.3 mm). The modifications were therefore tailored for self-calibrations of fainter sources. We describe the detailed processes below.
@@ 3.2 Imaging Processes
For the imaging process, we utilized the tclean task in CASA with the multiscale clean option, employing scales of [0, 4, 12, 24, 48] pixels for Band 6 images and [0, 4] for Band 7 images to effectively capture structures at different sizes. Band 6 images were produced with a resolution of 1600$\times$1600 pixels and a pixel size of $\farcs$025, typical for achieving a CLEAN beam of $\farcs$35. For Band 7 images, we used 864$\times$864 pixels and a pixel size of $\farcs$04 to accommodate a larger beam size of $\farcs$7.
To enhance the fidelity of CLEAN component locations, a binary image plane mask was employed in CLEAN processes. Given that most Lupus and Upper Sco sources exhibit line emissions following a Keplerian rotation pattern, we constructed “cleaning” Keplerian masks for these sources. All parameters for these Keplerian masks, such as stellar mass, disk inclination angle, and position angle measured from continuum visibility fitting (see Table 2.1 and Vioque et al. 2025, in press) can be found in Deng et al. (2025, in press); Agurto-Gangas et al. (2025, in press). Masks were constructed using the keplerian_mask function [^5] https://github.com/richteague/keplerian_mask. A conservative maximum radius (Rmax) was used for each disk, based on eye evaluation to encompass the observed extent of 12CO line emission. For imaging the 12CO line of Lupus and Upper sources, we use uvrange =‘$>$50klambda’ to remove cloud contamination at scale$>$4″ (except for Lupus 10, which has a large disk). The choice of 4″ is based on the size of 12CO images in all Class II sources except for Lupus 10.
For the Ophiuchus sources, where line emissions are often contaminated by the envelope and/or outflows, we utilized the automask option in tclean, as detailed in Ruiz-Rodriguez et al. (2025, in press). No baseline filtering was used for Ophiuchus sources.
@@
Figure 3: 1.3 mm (Left panels) and 12CO (2-1) moment zero images (Right panels) of the AGE-PRO sample. The orders are from the brightest 1.3 mm continuum source to the faintest one within each star-forming region. Images are 2″$\times$2″and colors are peak normalized and shown with a linear color stretch. The exception is Lupus 10 (V1094 Sco) which is shown in a 5″$\times$5″image and its continuum image is shown with an asinh stretch to accentuate fainter details. The difference in image size between Lupus 10 and the other sources is visualized in cyan in its continuum image. The beam sizes are shown in a circle on the lower-left of each panel and a 50 au size bar is shown in the upper right corner of each panel.

### v2

@@ Abstract
We present the ALMA Survey of Gas Evolution of PROtoplanetary Disks (AGE-PRO), a Large Program of the Atacama Large Millimeter/submillimeter Array (ALMA). AGE-PRO aims to systematically trace the evolution of gas disk mass and size throughout the lifetime of protoplanetary disks. It uses a carefully selected sample of 30 disks around M3-K6 stars in three nearby star-forming regions: Ophiuchus (0.5-1 Myr), Lupus (1-3 Myr), and Upper Sco (2-6 Myr). Assuming the three regions had similar initial conditions and evolutionary paths, we find the median gas disk mass appears to decrease with age. Ophiuchus disks have the highest median gas mass (6 MJup), while the Lupus and Upper Sco disks have significantly lower median masses (0.68 and 0.44 MJup, respectively). Notably, the gas and dust disk masses appear to evolve on different timescales. This is evidenced by the median gas-to-dust mass ratio, which decreases from 122 in the youngest disks (<<1 Myr) to 46 in Lupus disks, and then increases to 120 in the Upper Sco disks. The median gas disk sizes range between 74-110 au, suggesting that typical gas disks are much smaller than those of well-studied, massive disks. Population synthesis models suggest that magneto-hydrodynamic wind-driven accretion can reproduce median disk properties across all three regions, when assuming compact disks with a declining magnetic field over time. In contrast, turbulent-driven models overestimate gas masses of >>1 Myr disks by an order of magnitude. Here we discuss the program’s motivation, survey design, sample selection, observation and data calibration processes, and highlight the initial results.
@@ 1 Introduction
Planets form in gas- and dust-rich circumstellar disks around young stars (e.g., see reviews by Williams & Cieza, 2011; Andrews, 2020; Öberg et al., 2023). Most of these disks last for up to 3-10 Myr before the gas and dust dissipate, which sets the timescale of giant planet formation (e.g., Fedele et al., 2010; Ribas et al., 2014). The structure and evolution of these disks profoundly affect every step of planet formation: from the initial grain growth, the formation of planetesimals, the accretion of planetary atmospheres, to the migration of planets (e.g., Benz et al., 2014; Drazkowska et al., 2023; Krijt et al., 2023).
@@ 2.1 Sample Selection
Figure 1: Top: Hertzsprung-Russell diagram of the AGE-PRO sample. The pre-main-sequence evolutionary tracks and isochrones are from Feiden (2016); Baraffe et al. (2015). Ophiuchus sources do not have well-constrained stellar luminosity due to high extinction. We instead plot their luminosity ranges based on statistical ages of Class I and Flat spectrum sources (0.5-1 Myr) (Evans et al., 2009). Bottom: 1.3mm-continuum luminosity ($L_{\rm mm}$) distributions (using AGE-PRO continuum fluxes and normalized to 140 pc) of the selected sample in each region. The sample covers the entire $L_{\rm mm}$ range of M3-K6 sources in each star-forming region (see detailed comparisons in Ruiz-Rodriguez et al., 2025; Deng et al., 2025; Agurto-Gangas et al., 2025).
@@
References. — All distances are adopted from Bailer-Jones et al. (2021). In Col. (13), the references for Class, Spectral type, $T_{\rm eff}$, $L_{\ast}$}, Av, and accretion rates: 1 = Evans et al. (2009), 2 = Cieza et al. (2019), 3 = Ruiz-Rodriguez et al. (2025), 4 = Alcalá et al. (2014), 5 = Alcalá et al. (2017), 6 = Deng et al. (2025), 7 = Luhman (2022), 8 = Manara et al. (2020), 9 = Carpenter et al. 2024, submitted, 10 = Manara et al. (2023), 11 = Fang et al. (2023), 12 = Agurto-Gangas et al. (2025).
The AGE-PRO sample was selected with four main criteria: (1) sources with known stellar spectral type between M3-K6  [^1] Upper Sco 2 was classified as M2 in Preibisch et al. (2001) which we adopted in the AGE-PRO sample selection stage. But more recently Manara et al. (2020) reclassified the source as M4.5 with X-shooter spectra. , roughly corresponding to a stellar mass (M) of 0.3-0.8 M. The majority of stellar spectral types were adopted from literature and the rest from Ruiz-Rodriguez in prep. The literature stellar luminosities were updated with Gaia DR3 distances (see the list of references in Table 2.1). The focus on a narrow space of stellar mass is because disk properties are expected to depend on stellar mass (e.g., dust disk masses are known to be correlated with M; Pascucci et al. 2016; Andrews et al. 2018a). The stellar properties of the AGE-PRO sample are listed in Table 2.1. (2) sources without known companions or that are in wide-separation binaries (>>600 AU), as close binaries may evolve differently due to tidal interactions (e.g., Cuello et al., 2023). (3) For the Ophiuchus region, Class I/Flat spectrum sources were chosen to ensure the sub-sample is <<1 Myr old (Evans et al., 2009); for Lupus and Upper Sco regions, Class II sources were selected to exclude debris disks. (4) sources were selected from previous detections of mm continuum and CO line emission. These four criteria led to 10-34 viable candidates for each region. To ensure that our disks are representative of the selected M range, ten disks were selected from each region that covers the spread of continuum luminosities in the region (Ansdell et al. 2016; Barenfeld et al. 2016; Cieza et al. 2019; Carpenter et al. 2025). We excluded known edge-on disks (>>70 deg) as they have smaller emitting areas and thus have fainter line emission. Figure 1 shows the stellar properties of the AGE-PRO sample in the Hertzsprung-Russell diagram and their continuum luminosity distributions.
@@ 2.2 Ages of the AGE-PRO sample
Since the primary goal of the AGE-PRO program is to study the evolution of gas disks, constraining the ages of the sample is crucial. However, estimating the absolute ages of very young stars (<<10 Myr, all pre-main sequence PMS) is challenging due to the scarcity of direct age indicators and limited knowledge of star formation processes (see review by Soderblom et al. 2014). Typically, the ages of young PMSs are estimated by comparing observations with theoretical evolutionary tracks in either Hertzsprung-Russell Diagrams (HRDs) or color-magnitude diagrams (CMDs). If the stellar parameters and extinctions of individual stars are sufficiently constrained, ages can be determined fitting evolutionary tracks to each star, though these individual ages are often considered less reliable. Median ages are then adopted for the ages of clusters or star-forming regions. Another technique, often used when the extinctions and spectral types of individual stars are uncertain, is to derive a single age for an entire cluster by fitting all stars with one theoretical isochrone, assuming all clustered stars to be coeval.
@@ 2.3 Stellar masses of the AGE-PRO sample
In Figure 2 we compare the distribution of stellar masses across the three star-forming regions. The stellar masses are derived from the comparison with evolutionary tracks as described above. The bolometric temperature of our Ophiuchus sample indicates these are more evolved embedded disks, and therefore ,their current stellar masses should be close to their final masses. The median stellar masses of the three regions 0.45, 0.40, and 0.43 M for the Ophiuchus, Lupus, and Upper Sco region, respectively. The uncertainties of stellar masses from our estimations are typically 0.1-0.2 M, based on the uncertainties of the stellar effective temperature and luminosity (see details in Deng et al., 2025; Agurto-Gangas et al., 2025). However, these uncertainties do not include uncertainties of evolutionary tracks which are known to be model-dependent (Soderblom et al., 2014). Comparisons between stellar masses derived by evolutionary tracks with dynamical masses measured in binary systems suggest typical 30-80% uncertainties (Simon et al., 2019). Given these uncertainties on stellar masses, we consider the stellar masses in our sample are similar across the three regions and therefore suitable for comparison.
@@ 2.4.1 Lines to measure gas disk masses
But the conversion from CO gas mass to the total gas disk mass suffers from the high uncertainty of CO-to-H2 ratio ($x_{\rm CO}$) (Favre et al. 2013; Schwarz et al. 2016; Bergin & Williams 2017; Miotello et al. 2023). The uncertainty of $x_{\rm CO}$ likely increases over time due to chemical processes converting CO into more complex carbon carriers and dust growth carries CO ice from the disk atmosphere to the mid-plane (e.g, Schwarz et al., 2016; Bosman et al., 2018; Krijt et al., 2018, 2020). This expectation has been seen in some recent observations. Zhang et al. (2020); Bergner et al. (2020) compared CO isotopologue line fluxes with those predicted by thermo-chemical models and found that C/H and O/H elemental ratios of Class I/FS disks is consistent with the Interstellar Medium (ISM) ratio ($\sim$10-4). In contrast, they found that the C/H and O/H abundance of Class II disks needs to be reduced to $0.1-0.01\times$ ISM level. This is consistent with the timescales of the chemical processing of CO and the growth from ISM-sized dust particles to pebbles being $\sim$1 Myr (e.g., Schwarz et al., 2018; Bosman et al., 2018; Krijt et al., 2020). However, other works that include conversion of CO into CO2 ice do not require significant reductions in the C/H and O/H elemental abundances with respect to ISM values (Ruaud et al., 2022; Deng et al., 2023; Pascucci et al., 2023b; Ruaud & Gorti, 2024).
[…]
@@
Given these observations and theoretical expectations, the AGE-PRO program adopted different approaches to measure gas masses for sources younger and older than 1 Myr. For Ophiuchus sources (<<1 Myr), we used C17O (2-1) and assume ISM level $x_{\rm CO}$ of 10-4 to measure gas disk masses. Based on the dust masses and an ISM level of CO abundance, radiative transfer models predict that C18O (2-1) is optically thick and therefore the more rare C17O line (typical C18O/C17O abundance ratio of 3.6, Wilson 1999) is used for mass measurements (Miotello et al., 2016). For Class II disks, $x_{\rm CO}$ needs to be constrained for accurate gas measurements, and therefore 13CO, C18O (2-1), and N2H+ (3-2) lines were used to constrain $x_{\rm CO}$ and the gas disk masses.
@@ 3 Observations
The ALMA observations of the AGE-PRO program were conducted with the 12-meter array from 2021 December to 2022 August, with baseline lengths between 15 to 2917 m. Observational logs of each execution, including observational dates, array configurations, baseline length range, number of antennas, phase and flux calibrators, are shown in Ruiz-Rodriguez et al. (2025); Deng et al. (2025); Agurto-Gangas et al. (2025), for sources in Ophiuchus, Lupus, and Upper Sco, respectively.
These observations were carried out at ALMA band 6 and 7 (Ediss et al., 2004; Mahieu et al., 2012). Table 2 lists the correlator set-ups used for the AGE-PRO program. For Ophiuchus sources, two Band 6 setups were used: one for 12CO/13CO/C18JJ=2-1 lines, and the other one for C17JJ=2-1 line. The average on-source time for Ophiuchus targets was $\sim$30 minutes for the 12CO setup and $\sim$60 minutes for the C17O setup. For Lupus and Upper Sco sources, one Band 6 setup covered the CO/13CO/C18O/ JJ=2-1 lines and one Band 7 setup for the N2H+ JJ=3-2 line. The average on-source time of the Band 6 CO setup was $\sim$30 and 60 minutes for the Lupus and Upper Sco targets, respectively. The Band 7 N2H+ setup had 60 and 120  minutes of on-source time for the Lupus and Upper Sco targets, respectively. In addition to the key science lines, these spectral setups also covered various molecular lines for serendipitous discoveries, including H2CO, DCN, DCO+, N2D+, and CH3CN. One or two continuum bands with 1875 MHz bandwidth were included in all correlator setups for self-calibration. Because Upper Sco disks are generally the faintest in continuum emission among the AGE-PRO sample, we used a slightly different spectral setup in Band 6 for Upper Sco than for Lupus sources, doubling the continuum bandwidth to obtain higher sensitivity, which lowers the spectral resolution for some non-key science lines.
@@ 3.1 Pipeline calibration and self-calibration
After the standard pipeline calibration, we conducted self-calibration using CASA version 6.4 to improve the signal-to-noise (SNR) of the data. Self-calibration of the continuum visibilities was performed for sources with continuum SNR larger than 20. Our methodology was based on the self-cal procedures of the DSHARP and MAPS large programs (Andrews et al., 2018b; Czekala et al., 2021), with modifications tailored to the AGE-PRO program. Unlike the targets of the DSHARP and MAPS large programs which were bright in the millimeter continuum, many of our sources only have intermediate and faint continuum emission (<<10 mJy at 1.3 mm). The modifications were therefore tailored for self-calibrations of fainter sources. We describe the detailed processes below.
@@ 3.2 Imaging Processes
[…]
To enhance the fidelity of CLEAN component locations, a binary image plane mask was employed in CLEAN processes. Given that most Lupus and Upper Sco sources exhibit line emissions following a Keplerian rotation pattern, we constructed “cleaning” Keplerian masks for these sources. All parameters for these Keplerian masks, such as stellar mass, disk inclination angle, and position angle measured from continuum visibility fitting (see Table 2.1 and Vioque et al. 2025) can be found in Deng et al. (2025); Agurto-Gangas et al. (2025). Masks were constructed using the keplerian_mask function [^5] https://github.com/richteague/keplerian_mask. A conservative maximum radius (Rmax) was used for each disk, based on eye evaluation to encompass the observed extent of 12CO line emission. For imaging the 12CO line of Lupus and Upper sources, we use uvrange =‘>>50klambda’ to remove cloud contamination at scale>>4″ (except for Lupus 10, which has a large disk). The choice of 4″ is based on the size of 12CO images in all Class II sources except for Lupus 10.
For the Ophiuchus sources, where line emissions are often contaminated by the envelope and/or outflows, we utilized the automask option in tclean, as detailed in Ruiz-Rodriguez et al. (2025). No baseline filtering was used for Ophiuchus sources.
@@
Figure 3: 1.3 mm (Left panels) and 12CO (2-1) moment zero images (Right panels) of the AGE-PRO sample. The orders are from the brightest 1.3 mm continuum source to the faintest one within each star-forming region. Images are $\times2$″and colors are peak normalized and shown with a linear color stretch. The exception is Lupus 10 (V1094 Sco) which is shown in a $\times5$″image and its continuum image is shown with an asinh stretch to accentuate fainter details. The difference in image size between Lupus 10 and the other sources is visualized in cyan in its continuum image. The beam sizes are shown in a circle on the lower-left of each panel and a 50 au size bar is shown in the upper right corner of each panel.